# Multiple Instance Verification

### v1

@@ 6 Experiments
We conducted experiments on the three tasks illustrated in Figure 3, also described in Section 3.1, based on the publicly available raw data [38, 24, 33]. For each task, we conducted three rounds of experiments and report mean performance and standard errors. Within each round of experiments, by randomly sampling from disjoint subsets of the raw data (i.e., with non-overlapping query classes, cf., Section 3.1), we constructed a different set of train, validation, and test data respectively. Note that only the test dataset includes the key instances’ ID, for evaluating quality of the explanations. In all tables, the best performance is indicated in bold. Appendix B has details on the raw data, feature extractors [14, 31, 28, 19], and training process used in each task.
@@ 6.1 Results for QMNIST handwriting verification
We first consider classification performance and quality of the explanations using training datasets with a fixed sample size and without controlling bag sizes, and then go on to consider the effect of varying training sample size and (controlled) average bag sizes.
6.1.1 Classification performance and quality of the explanations
Table 1: QMNIST: classification performance (mean and standard errors), including two non-attention-based benchmarks (MI-Net and Bi-LSTM). The best performance is shown in bold.
{NiceTabular}
@l@  c@ c@ c@ c@ c@ AUROCAccuracyPrecisionRecallF1-score
Baseline0.708$\pm$0.0060.658$\pm$0.0030.647$\pm$0.0030.637$\pm$0.0030.637$\pm$0.003
GABMIL0.660$\pm$0.0170.623$\pm$0.0110.613$\pm$0.0090.613$\pm$0.0090.613$\pm$0.009
PMA0.641$\pm$0.0050.604$\pm$0.0100.600$\pm$0.0060.600$\pm$0.0060.597$\pm$0.009
MSA0.634$\pm$0.0080.596$\pm$0.0040.593$\pm$0.0070.597$\pm$0.0030.590$\pm$0.006
MI-Net 0.637$\pm$0.006 0.593$\pm$0.005 0.597$\pm$0.003 0.597$\pm$0.003 0.593$\pm$0.003
Bi-LSTM 0.630$\pm$0.005 0.594$\pm$0.010 0.590$\pm$0.006 0.590$\pm$0.006 0.587$\pm$0.009
CAP-VEMA(ours)0.736$\pm$0.0040.675$\pm$0.0070.663$\pm$0.0070.667$\pm$0.0090.667$\pm$0.009
CAP-DBA(ours)0.731$\pm$0.0110.675$\pm$0.0090.667$\pm$0.0070.667$\pm$0.0070.667$\pm$0.007
Table 2: Classification performance of CAP-VEMA model on a test QMNIST dataset broken down by digits
@@ 6.1.1 Classification performance and quality of the explanations
Table 3: QMNIST: evaluating quality of the explanations (mean and standard errors). The best performance is shown in bold.
{NiceTabular}
@l@cl@ Avg. i-AUROC Avg. i-AP
@@
Table 6.1.1 reports the classification performance of all methods, including the two non-attention-based ones (MI-Net and Bi-LSTM), on QMNIST using AUROC, accuracy, precision, recall, and F1-score. [^8] We use the macro average for precision, recall, and F1-score because either class may be of interest, e.g., the exemplar label 0 when models are used for anomaly detection, and the label 1 for verification purposes. It shows the baseline outperforms the benchmark models, while the CAP-based models improve on the baseline and benchmarks in all measures, with statistical significance (at $\geq 95\%$ confidence level) and by substantial margins.
@@
To better understand the exemplar-level verification performance in Table 6.1.1, it is worth noting that even though MNIST is generally regarded as a simple dataset, the MI verification task based on it can be highly challenging. To put it in context, for the well-known task of image classification to recognize MNIST digits, usually deemed as a “solved” problem in the literature, a ResNet-18 model can typically achieve above 95% classification accuracy [13]. In contrast, the accuracy of all models in Table 6.1.1, using ResNet-18 as a feature extractor, is below 70% on the MI verification task. In addition, our experimental design, constructing each exemplar using the same digit, enables us to break down performance by digits, as shown in Table 6.1.1, which facilitates understanding and disentangling relevant factors. Table 6.1.1 has performance per digit evaluated using the CAP-VEMA model based on one representative test dataset of this task.
@@
We offer some intuition about why the MNIST MI-verification task can be difficult and draw some observations from Table 6.1.1. Intuitively, the high achievable accuracy, and thus easiness, of the digit recognition task demonstrates that the writing styles of different writers are generally consistent—otherwise inconsistent or ambiguous writing patterns would have made it much more difficult to recognize, or differentiate between, digits [25, 26]. Nonetheless, such consistency of writing styles is in fact a major source of difficulty for a verification task, because the goal of verification is to tell apart different writers’ handwriting based on variations in writing style. [^9] In an extreme case, had all writers’ writing style been identical, it would have made the digit recognition task trivial, but made the verification task impossible—that is, no different from random guess—even for humans. MI verification adds further challenge by introducing bags of different writers’ handwriting. This exacerbates the issue of less variation in writing styles since only one confounding writer in a bag can result in the signal being overwhelmed by the noise. Table 6.1.1 demonstrates some supporting evidence for this intuition: the digits usually written in a uniform pattern, like “1” and “8”, are more difficult for MI verification, leading to higher false positive/negative rates and thus worse accuracy. Conversely, digits that can be naturally written in diverse ways, such as “2” and “7”, can be more accurately verified.
@@
To enable evaluation of explanation quality, Table 3 shows average i-AUROC and i-AP for the baseline, two benchmarks, and our models. CAP-based models outperform the other three by even larger margins than in the case of classification performance, indicating a superior ability to identify key instances. Notably, the two benchmarks produce significantly worse explanation quality than the baseline, supporting the hypothesis that failure of their attention mechanism to incorporate the query diminishes their ability to accurately model verification tasks of this type.
@@ 6.1.2 Effects of training sample size and bag size
(b) Quality of the explanations: average i-AUROC
@@
(d) Quality of the explanations: average i-AP
Figure 5: Test AUROC (4(a)), accuracy (4(c)), average i-AUROC (4(b)), average i-AP (4(d)) for QMNIST when varying numbers of training exemplars, from 8,000 to 40,000, with an average of 10 instances per target bag.
Similar to [17], we vary the number of training exemplars and the number of instances per target bag in the training data, and study the effect on model performance. More specifically, we consider mean bag sizes of 10, 20, and 50, with variance 2, 4, and 10, respectively. For mean bag size 10, Figures 5 shows classification performance (AUROC, accuracy) and quality of explanations (average i-AUOROC, average i-AP), obtained by varying the number of training exemplars from 8,000 to 40,000. The results for mean bag sizes 20 and 50 can be found in Appendix C, Figures 7 and 8, and are very similar.
Based on these results, it is clear that CAP-based models perform substantially better than the other models, for small and large numbers of training examples. The improvement in quality of the explanations (cf. Figure 4(b) and 4(d)) is consistent with the hypothesis that this is the main driver for the improvement in classification performance. For average target bag sizes of 20 and 50 instances, similar to the case of bag size 10, our models still significantly outperform other models by substantial margins, according to all criteria.
We also note that all models’ performance deteriorates when the number of instances per target bag increases. A deeper inspection of the data explains the reason. In this task, the source of key instances, i.e., the handwriting of a specific digit from a writer, is limited in the raw data. Because we draw random samples without replacement when generating a bag, key instances in a bag become relatively scarce when we increase the total number of instances in the bag. This can be clearly seen in the second column of Table 4, which presents the mean/median/maximum statistics for the number of key instances per target bag in the training data.
Average number of instances per bag Number of key instances (Number of key instancesNumber of $\Big{(}\frac{\text{Number of key instances}}{\text{Number of instances}}\Big{)}$
@@
Table 4 shows that when the average bag size increases, the number of key instances per target bag remains stable. Consequently, enlarging the bag size entails adding more non-key instances, or “noise”, to the bag. In other words, the ratio of “Number of key instances / Total number of instances” per bag, akin to a signal-to-noise ratio, decreases with increased bag sizes, see the third column of Table 4. Therefore, in this task, the “number of instances per target bag” effectively becomes a proxy for “noise-to-signal” ratio; the results show that higher noise-to-signal ratio in the training data creates a more difficult verification task, which causes performance deterioration for all models. From Figures 7 and 8, we also observe that the performance deterioration of our models tends to be less sensitive to the increase of bag size, indicating that our models are more resilient to higher noise-to-signal ratios.
@@ 6.2 Results for signature verification against multiple anchors
{NiceTabular}
@l@  c@ c@ c@ c@ c@ AUROC Accuracy Precision Recall F1-score
Baseline 0.685$\pm$0.014 0.594$\pm$0.008 0.697$\pm$0.019 0.593$\pm$0.009 0.533$\pm$0.015
GABMIL 0.688$\pm$0.035 0.547$\pm$0.018 0.697$\pm$0.007 0.547$\pm$0.018 0.440$\pm$0.036
@@
MSA 0.690$\pm$0.011 0.619$\pm$0.002 0.633$\pm$0.003 0.620$\pm$0.000 0.610$\pm$0.000
MI-Net 0.703$\pm$0.039 0.643$\pm$0.016 0.673$\pm$0.007 0.643$\pm$0.017 0.623$\pm$0.022
Bi-LSTM 0.733$\pm$0.027 0.653$\pm$0.019 0.663$\pm$0.024 0.653$\pm$0.020 0.647$\pm$0.018
CAP-VEMA(ours) 0.819$\pm$0.006 0.673$\pm$0.027 0.733$\pm$0.003 0.673$\pm$0.029 0.647$\pm$0.041
CAP-DBA(ours) 0.814$\pm$0.006 0.674$\pm$0.018 0.733$\pm$0.009 0.673$\pm$0.020 0.653$\pm$0.029
Table 6: Signature verification against multiple anchors: evaluating the explanations (mean and standard errors). The best performance is shown in bold.
{NiceTabular}
@l@cc@ Avg. i-AUROC Avg. i-AP
@@
Figure 6: More examples to illustrate quantitative evaluations of explanations vs. classification. The column “Predicted Probability” shows a model’s predicted probability of the exemplar label. If it is $\geq 0.5$, this exemplar is classified as 1, otherwise 0. Ground-truth of a key instance is highlighted by a red rectangle. Signature images are blurred for data use compliance.
@@
The main results for signature verification are in Tables 6.2 and 6, respectively. The CAP-based models consistently exceed the performance of the baseline and benchmarks by sizable margins, according to both predictive performance and explanation quality. In particular, they provide near-perfect explanations, which supports the hypothesis that their superior classification performance is due to their ability to correctly identify key instances.
@@
Overall, the pattern of performance in Tables 6.2 and 6 is similar to that in the first task, with two exceptions. One, our models’ explanation quality is close to perfect whilst their classification accuracy is substantially below 100%. Two, the PMA benchmark appears to perform notably better than the baseline, although with higher standard errors. Considering the first point, as explained in Section 3.1, there are two types of negative exemplars in this task: one has an irrelevant query that is completely different from any anchor; another includes a query that is a professional forgery similar to an (authentic) anchor, sometimes looking even more like the anchors than some truly authentic queries. Because the training data do not disclose what type of query yields a negative exemplar (“irrelevant” or “forgery”)—to mimic the real-world scenarios where the reasons for rejecting a signature are unavailable—solving this task requires essentially a two-step verification process: (1) to identify the correct anchors from a target bag, if they exist; and (2) to verify whether the query is authentic or a forgery when compared to relevant anchors. This process is the same even if humans perform the verification—the first step may be relatively easy while the second can be highly difficult. The near-perfect explanations demonstrate our methods’ ability to perform the first step well while the relatively low accuracy reflects the difficulty of distinguishing between authentic signatures and forgeries in the second step.
@@
Considering PMA’s relatively high classification accuracy, inspection of its explanation quality reveals that correct classifications are often obtained for the wrong reasons—its explanation quality is far worse than that of the baseline. This implies that even though some classifications may be “correct”, the incorrect key instances in the target bag are chosen when making those classifications. This faulty behaviour is enabled by the fact that PMA uses a trainable seed vector, rather than the query’s feature vector, for MI pooling of the target bag. Intuitively, using incorrect reasoning to obtain classifications implies high variance (standard errors) in classification performance and this is confirmed in Table 6.2. The discrepancy in average classification performance and explanation quality highlights the importance of assessing both criteria together.
@@
Examples of inconsistency in the accuracy of classifications and quality of the explanations are shown in Figures 5(a) and 5(b). In Figure 5(a), there is one key instance: the third instance. Our models correctly identify this key instance with the highest attention score (which is much higher than the second-highest score) and also make the correct classification. The highest attention score of the baseline and GABMIL benchmark does not identify the correct key instance, although the GABMIL’s second-highest score does. Note that both i-AUROC and i-AP correctly reflect GABMIL’s better quality of the explanations than the baseline’s. In both cases, they make an incorrect classification, as expected. The PMA benchmark has the poorest quality of explanations, as expected, because its lowest attention score corresponds to the actual key instance. Surprisingly, it makes the correct classification with a high predicted probability.
@@
For the exemplar in Figure 5(b), there are two key instances. Our models correctly identify them with the two highest attention scores. Nonetheless, the DBA method makes an incorrect classification, despite the good quality of the explanations. Likewise, the baseline correctly identifies one key instance yet still makes an incorrect classification. Both benchmarks, GABMIL and PMA, are unable to identify the correct key instances, with the two lowest attention scores corresponding to the ground truth. Whilst GABMIL makes an incorrect classification as expected, interestingly PMA still makes the correct classification.

### v2

@@ 6 Experiments
We conducted experiments on the three tasks illustrated in Figure 4, also described in Section 3.1, based on the publicly available raw data in (Yadav and Bottou (2019); Liwicki et al. (2011); Thorne et al. (2018)). For each task, we conducted three rounds of experiments and report mean performance and standard errors. Within each round of experiments, by randomly sampling from disjoint subsets of the raw data (i.e., with non-overlapping query classes, cf., Section 3.1), we constructed a different set of train, validation, and test data, respectively. Note that only the test dataset includes the key instances’ ID, for evaluating key instance detection performance. In all tables, the best performance is indicated in bold. Appendix B has details on the raw data, feature extractors (from He et al. (2016); Tan and Le (2021); Reimers and Gurevych (2019); Kanavati and Tsuneki (2021)), and training process used in each task.
For our proposed approach, i.e., CAP, we use “CAP-VEMA” and “CAP-DBA” to refer to the models with VEMA and “DBA with $L_{1}$-distance” attention functions, respectively. The exception is our study on the QMNIST task, where we evaluated DBA with both $L_{1}$- and $L_{2}$-distance, denoted as CAP-DBA-$L_{1}$ and CAP-DBA-$L_{2}$ in Tables 2 and 4.101010We thank an anonymous reviewer for suggesting the experiments with $L_{2}$-distance.
@@ 6.1 Results for QMNIST handwriting verification
We first consider performance of classification and key instance detection using datasets with a fixed training sample size and without controlling bag sizes, and then go on to consider the effect of varying training sample size and (controlled) average bag sizes.
6.1.1 Performance of classification and key instance detection
Table 2: QMNIST: classification performance (mean and standard errors), including non-attention-based benchmarks (MI-Net, Bi-LSTM). The best performance is shown in bold.
AUROC Accuracy Precision Recall F1-score
[…]
Table 3: Classification performance of CAP-VEMA model on test QMNIST dataset broken down by digits
@@ 6.1.1 Performance of classification and key instance detection
Table 4: QMNIST: key instance detection performance (mean and standard errors). The best performance is shown in bold.
Avg. i-AUROC Avg. i-AP
@@
Table 2 reports the classification performance of all methods, including the two non-attention-based ones (MI-Net and Bi-LSTM), on QMNIST using AUROC, accuracy, precision, recall, and F1-score. [^11] We use the macro average for precision, recall, F1-score because either class may be of interest, e.g., the exemplar label 0 when models are used for anomaly detection, and the label 1 for verification purposes. It shows the baseline outperforms the benchmark models, while the CAP-based models improve on all measures by substantial margins when compared with the baseline and benchmarks. Interestingly, the results using $L_{2}$-distance, at least on the QMNIST datasets, are slightly better than those using DBA-$L_{1}$ and VEMA.
@@
To better understand the exemplar-level verification performance in Table 2, it is worth noting that even though MNIST is generally regarded as a simple dataset, the MI verification (MIV) task based on it can be highly challenging. To put it in context, for the well-known task of image classification to recognize MNIST digits, usually deemed a “solved” problem in the literature, a ResNet-18 model can typically achieve above 95% classification accuracy (Ghosh et al. (2023)). In contrast, the accuracy of all models in Table 2, using ResNet-18 as a feature extractor, is below 70% on the MIV task. In addition, our experimental design, constructing each exemplar using the same digit, enables us to break down performance by digits, as shown in Table 3, which facilitates understanding and disentangling relevant factors. Table 3 has performance per digit evaluated using the CAP-VEMA model based on one representative test dataset of this task.
@@
We offer some intuition about why the MNIST MIV task can be difficult and draw some observations from Table 3. Intuitively, the high achievable accuracy, and thus easiness, of the digit recognition task demonstrates that the writing styles of different writers are generally consistent—otherwise inconsistent or ambiguous writing patterns would have made it much more difficult to recognize or differentiate between digits (Mukhoti et al. (2021); Park et al. (2023)). Nonetheless, such consistency of writing styles is in fact a major source of difficulty for a verification task, because the goal of verification is to tell apart different writers’ handwriting based on variations in writing style. [^12] In an extreme case, had all writers’ writing style been identical, it would have made the digit recognition task trivial, but made the verification task impossible—that is, close to random guess—even for humans. MIV adds further challenge by introducing bags of different writers’ handwriting. This exacerbates the issue of less variation in writing styles since only one confounding writer in a bag can result in the signal being overwhelmed by the noise. Table 3 contains some supporting evidence for this intuition: the digits usually written in a uniform pattern, like “1” and “8”, are more difficult for MIV, leading to higher false positive/negative rates and thus worse accuracy. Conversely, digits that can be naturally written in diverse ways, such as “2” and “7”, can be more accurately verified.
@@
To enable evaluation of key instance detection performance, Table 4 shows average i-AUROC and i-AP for the baseline, two benchmarks, and our models. CAP-based models outperform the other three by even larger margins than in the case of classification performance, indicating a superior ability to identify key instances based on the attention scores. Notably, the two benchmarks produce significantly worse performance than the baseline, supporting the hypothesis that failure of their attention mechanism to incorporate the query diminishes their ability to effectively tackle verification tasks of this type.
@@ 6.1.2 Effects of training sample size and bag size
(b) Key instance detection performance: average i-AUROC
@@
(d) Key instance detection performance: average i-AP
Figure 6: Test AUROC (6(a)), accuracy (6(c)), average i-AUROC (6(b)), average i-AP (6(d)) for QMNIST when varying numbers of training exemplars, from 8,000 to 40,000, with an average of 10 instances per target bag.
Similar to Ilse et al. (2018), we vary the number of exemplars (training sample size) and the number of instances per target bag (bag size) in the training data, and study the effect on model performance. More specifically, we consider mean bag sizes of 10, 20, and 50, with variance 2, 4, and 10, respectively. For mean bag size 10, Figure 6 shows performance of classification (AUROC, accuracy) and key instance detection (average i-AUOROC, average i-AP), obtained by varying the training sample size from 8,000 to 40,000. The results for mean bag sizes 20 and 50 can be found in Figures 89 of Appendix C and are very similar. In this experiment, we only consider $L_{1}$-distance for the DBA attention function.
Based on these results, it is clear that CAP-based models perform substantially better than the other models, for small and large training sample sizes. The improvement in key instance detection performance (cf. Figure 6(b) and 6(d)) is consistent with the hypothesis that this is the main driver for the improvement in classification performance. For average target bag sizes of 20 and 50 instances, similar to the case of bag size 10, our models still significantly outperform the other models by substantial margins, according to all criteria.
Average number of instances per bag Number of key instances $\Big{(}\frac{\text{Number of key instances}}{\text{Number of instances}}\Big{)}$
@@
We also observe that all models’ performance worsens when the bag size increases. A deeper inspection of the data reveals the reason. In this task, the source of key instances, i.e., a set of handwritten examples of a specific digit from a writer, is limited in the raw data. Because we draw random samples, without replacement and by conditioning on a digit, to generate a bag, key instances in a bag become relatively scarce when we increase the bags size. This can be clearly seen in the second, third, and fourth column of Table 5, which present the mean, median, and maximum, respectively, for the number of key instances per target bag in the training data: when the average bag size increases, the number of key instances per bag remains stable. Consequently, enlarging the bag size entails adding more non-key instances, or “noise”, to the bag. In other words, the ratio “Number of key instances / Total number of instances” per bag, [^13] This ratio stands for the proportion of key instances in each bag, sometimes also known as “witness rate” (Carbonneau et al. (2018); Early et al. (2022)). representing a form of signal-to-noise ratio, decreases with increased bag size, see the rightmost three columns of Table 5. Therefore, in this task, the bag size effectively becomes a proxy for the signal-to-noise ratio. When complicated by low inter-class variation relative to intra-class variation in QMNIST, illustrated in Section 6.1.1, a low signal-to-noise ratio in the training data creates a more difficult verification task for any model or even humans, not just for the models considered here. Empirically, comparing Figures 8 and 9 to Figure 6, the greater difficulty due to larger bag size manifests itself in a drop in performance across all models. Similar observations, that a lower signal-to-noise ratio (i.e., witness rate) correlates with increased difficulty of tasks and weaker performance of all models, can also be found in the existing MIL literature (e.g., Early et al. (2022)).
Importantly, we note that the performance of the CAP-based models tends to be less sensitive to increases in bag size, indicating that these models are more resilient to lower signal-to-noise ratios. For example, when the average bag size increases from 10 to 20 (cf. Figure 8), these models’ performance varies little while that of the other models deteriorates substantially. For an average bag size of 50 (cf. Figure 9), performance of our models decreases only slightly or moderately, while that of other models is noticeably reduced.
@@ 6.2 Results for signature verification against multiple anchors
AUROC Accuracy Precision Recall F1-score
@@
[…]
Table 7: Signature verification against multiple anchors: key instance detection performance (mean and standard errors). The best performance is shown in bold.
Avg. i-AUROC Avg. i-AP
@@
Figure 7: More examples to illustrate quantitative evaluations of key instance detection vs. classification. The column “Predicted Probability” shows a model’s predicted probability of the exemplar label. If it is $\geq 0.5$, this exemplar is classified as 1, otherwise 0. A ground-truth key instance is highlighted by a red rectangle. Signature images are blurred for data use compliance.
@@
The main results for signature verification are in Tables 6 and 7, respectively. The CAP-based models consistently exceed the performance of the baseline and benchmarks by sizable margins, according to performance of both classification and key instance detection. In particular, they provide near-perfect key instance detection, which supports the hypothesis that their better classification performance is due to their superior ability to correctly identify key instances.
@@
Overall, the pattern of performance in Tables 6 and 7 is similar to that in the first task, with two exceptions. One, our models’ key instance detection performance is close to perfect whilst their classification accuracy is substantially below 100%. Two, the PMA benchmark appears to perform notably better than the baseline, although with higher standard errors. Considering the first point, as explained in Section 3.1, there are two types of negative exemplars in this task: one has an irrelevant query that is completely different from any anchor; another includes a query that is a professional forgery similar to an (authentic) anchor, sometimes looking even more like the anchors than some truly authentic queries. Because the training data do not disclose what type of query yields a negative exemplar (“irrelevant” or “forgery”)—to mimic the real-world scenarios where the reasons for rejecting a signature are unavailable—solving this task requires essentially a two-step verification process: (1) to identify the correct anchors from a target bag, if they exist; and (2) to verify whether the query is authentic or a forgery when compared to relevant anchors. This process is the same even if humans perform the verification—the first step may be relatively easy while the second can be highly difficult. The near-perfect key instance detection performance demonstrates our methods’ ability to perform the first step well while the relatively low accuracy reflects the difficulty of distinguishing between authentic signatures and forgeries in the second step.
@@
Considering PMA’s relatively high classification accuracy, inspection of its key instance detection performance reveals that “correct” classifications are often obtained for the wrong reasons—its key instance detection performance is far worse than that of the baseline. This implies that even though some classifications may be “correct”, the incorrect key instances in the target bag are chosen when making those classifications. This faulty behaviour is enabled by the fact that PMA uses a trainable seed vector, rather than the query’s feature vector, for MI pooling of the target bag. Intuitively, using incorrect reasoning to obtain classifications implies high variance (standard errors) in classification performance and this is confirmed in Table 6. The discrepancy between performance of classification and key instance detection, highlights the importance of assessing both criteria together.
@@
We show some concrete examples of inconsistency in the accuracy of classifications and key instance detection in Figures 7(a) and 7(b). In Figure 7(a), there is one key instance: the third instance. Our models correctly identify this key instance with the highest attention score (which is much higher than the second-highest score) and also make the correct classification. The highest attention score of the baseline and GABMIL benchmark does not identify the correct key instance, although GABMIL’s second-highest score does. Note that both i-AUROC and i-AP correctly reflect their worse ability to detect the key instance compared to our models, and as expected, they make an incorrect classification. The PMA benchmark has the poorest key instance detection performance, because its lowest attention score corresponds to the actual key instance. Surprisingly, it makes the correct classification with high probability.
@@
For the exemplar in Figure 7(b), there are two key instances. Our models correctly identify them with the two highest attention scores. Nonetheless, the DBA method makes an incorrect classification, despite the accurate detection of the key instances. Likewise, the baseline correctly identifies one key instance yet still makes an incorrect classification. Both GABMIL and PMA are unable to identify the correct key instances, with the two lowest attention scores corresponding to the true key instances. Interestingly, GABMIL makes an incorrect classification, but PMA still makes the correct classification.